Delta’s stock has 30% chance of producing a 15% return, a 40% chance of producing a 10% return, and a 30% chance of producing a 2% return. What is the firm's expected rate of return?
1
Expert's answer
2020-06-25T18:48:38-0400


according to the formula expected portfolio return:


"0.3\times15+0.4\times10+0.3\times2=4.5+4+0.6=9.1 \%" return.